OpenStudy (anonymous):

What does it mean to graph the function: f(x) = 2(x + 3/2)^2 - 3, plot the vertex, y-intercept and axis of symmetry. Also find some more values of f(x)? Please help!

OpenStudy (anonymous):

On my textbook it shows: x f(x) -0.5 f(-1)=2(-0.5 + 1.5)^2 - 3 = -1 - 1 f(-1) = 2(-1 + 1.5)^2 - 3 = -2.5 -2 -2.5 (same as F9-1) since the parabola is symmetric about x = -1.5) -2.5 -1 (same as f (-0.5) since the parabola is symmetric about x = -1.5)

OpenStudy (anonymous):

I don't understand..